Sum of the digits of the number (104n2+8+1)2 , where n is a positive integer is ?

Grade:10

1 Answers

Sourabh Singh IIT Patna
7 years ago
Hii

Be it any number just calculate the total number of zeros that will precede the unit digit in the number

eg 101*101 = 10201
1001 *1001 =1002001
10001*10001 = 100020001

Hence in any case the sum of digigts is always going to be 4